A stock just paid a dividend of $2.24. The dividend is expected to grow at 23.88% for three years and then grow at 3.32% therRound to 2 decimal places and possibly show work.
Answer #1

D1=(2.24*1.2388)=2.774912

D2=(2.774912*1.2388)=3.43756099

D3=(3.43756099*1.2388)=4.25845055

Value after year 3=(D3*Growth rate)/(Required rate-Growth rate)

=(4.25845055*1.0332)/(0.112-0.0332)

=55.8354202

Hence value of stock=Future dividend and value*Present value of discounting factor(rate%,time period)

=2.774912/1.112+3.43756099/1.112^2+4.25845055/1.112^3+55.8354202/1.112^3

=$48.98(Approx).